# Learning Reward Machines from Partially Observed Policies

### v2

@@ 5.4 Real-world Mice Navigation
We also applied our learning framework to the trajectories of real mice navigating (in the dark) in a 127-node labyrinth maze (Rosenberg et al., 2021) shown in Figure 5. Each node, labeled with a number in Figure 5(a), represents a state in the MDP. The mouse can select from 4 actions: $\{\mathrm{stay},\mathrm{right},\mathrm{left},\mathrm{reverse}\}$. State $116$ (middle-right) contains a water resource and is labeled $\mathbf{w}$, state 00 (center) is the home state labeled $\mathbf{h}$, and all other states are labeled $\mathbf{i}$. Two cohorts of 10 mice moved freely in the maze for 7 hours, with one cohort being water-restricted and the other was not. A sample water-restricted mouse trajectory is shown in Figure 5(b). Difference in water restriction condition resulted in different animal behavior between these two cohorts. For the purposes of our study, we only considered the water-restricted mice. We used the same dataset of trajectories from (Ashwood et al., 2022), which is comprised of 200 mouse trajectories, given as state-action pairs of length $22$ each. We set our probability threshold at $\alpha=0.001$. With $u_{\mathrm{max}}=2$, our algorithm learns the unique reward machine shown in Figure 5(c). This is consistent with the seen behaviors of water-restricted mice which first try to reach the water port and hydrate, after which their behaviors switch to exploring the maze or going back the home state (Zhu et al., 2024; Ashwood et al., 2022). More details can be found in Appendix B.3.
@@ A.1 Proof of Proposition 1
More precisely, $\delta_{\mathbf{u}}(u_{I},\sigma)=\delta_{\mathbf{u}}(u_{I},\sigma^{\prime})$. Due to the contrapositive of Lemma 1, we have a contradiction. Similarly, due to the full specification condition of Section 3.1.2 and (10), the support of the two prefix policies will be the same by construction. ∎
@@ B.1 Tabular BlockWorld MDP
2Construct product MDP $\mathcal{M}_{\mathrm{Prod}}=\bigl{(}\mathcal{S}^{\prime},\;\mathcal{A}^{\prime},\;\mathcal{P}^{\prime},\;\mu_{0}^{\prime},\;\gamma^{\prime},\;r^{\prime}\bigr{)}$ as in Section 2.2.
@@ 13       $\pi^{*}(a\mid s,u)\;\leftarrow\;\exp\!\bigl{(}Q_{k}(s,u,a)-V_{k}(s,u)\bigr{)}$
Input: MDP $\mathcal{M}$, reward machine $\mathcal{R}$, true product policy $\pi^{*}(a\mid s,u)$, # trajectories $N$, trajectory length $H$
@@ 13             $s\leftarrow s^{\prime}$;
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ SAT solutions
1000 20 84 0.05 24
3000 20 226 0.05 12
5000 20 304 0.05 4
10000 20 456 0.05 4
0.1M 20 1965 0.05 4
1M 20 6242 0.05 2
@@
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ SAT solutions
100 20 24 0.05 40
200 20 37 0.05 8
500 20 77 0.05 8
1000 20 130 0.05 4
0.1M 20 1680 0.05 4
1M 20 4771 0.05 4
@@
B.2 2D-Continuous Robotics Reacher Arm
@@ B.2.2 Discretizing the state space and sampling the policy
The MDP state space is generated by discretizing the shoulder angle $\theta_{1}$, the elbow angle $\theta_{2}$, and the angular velocities $\dot{\theta}_{1},\dot{\theta}_{2}$. The angles are discritized uniformly in $[-\pi,\pi]$ with a bin size of $10^{\circ}\approx 0.17$ radians. The angular velocities are discretized uniformly in $[-14,14]$ with a bin size of $0.25$ rad/s. This leads to $\approx 17.4M$ states in the MDP.
@@ B.3 Real-world Mice Navigation
For recovering the reward function, we first need to construct the product policy. However, real-world data generally do not perfectly fit the assumed mathematical models, i.e., the max entropy assumption. Since most of the states in the MDP are unvisited, and many action are not sampled at most of the states, the product policy is predominantly sparse ($\sim 92\%$ of the action probabilities are 00). We preprocess the policy by clipping its minimum value to $0.05$. This allows us to use the results from [Shehab et al., 2024] to compute the reward function, as we requires computing the $\log$ of the policy. The reward space is constrained to satisfy the feature-decomposition which the reward machine provides. However, there is generally no guarantee that we can find such a featurized reward function that can perfectly induce any product policy. With the shown reward in Figure 5(c), the norm difference between the induced policy and the true policy is $2.66$ (the norm difference with a uniformly random policy is $2.97$). By increasing $u_{\mathrm{max}}$ to $3$, one of the recovered reward machines is shown in Figure 9, which yields a slightly larger policy difference norm of $2.71$.

### v3

@@ 5.4 Real-world Mice Navigation
We also applied our learning framework to the trajectories of real mice navigating (in the dark) in a 127-node labyrinth maze (Rosenberg et al., 2021) shown in Figure 5. Each node, labeled with a number in Figure 5(a), represents a state in the MDP. The mouse can select from 4 actions: $\{\mathrm{stay},\mathrm{right},\mathrm{left},\mathrm{reverse}\}$. State $116$ (middle-right) contains a water resource and is labeled $\mathbf{w}$, state 0 (center) is the home state labeled $\mathbf{h}$, and all other states are labeled $\mathbf{i}$. Two cohorts of 10 mice moved freely in the maze for 7 hours, with one cohort being water-restricted and the other was not. A sample water-restricted mouse trajectory is shown in Figure 5(b). Difference in water restriction condition resulted in different animal behavior between these two cohorts. For the purposes of our study, we only considered the water-restricted mice. We used the same dataset of trajectories from (Ashwood et al., 2022), which is comprised of 200 mouse trajectories, given as state-action pairs of length $22$ each. We set our probability threshold at $\alpha=0.001$. With $u_{\mathrm{max}}=2$, our algorithm learns the unique reward machine shown in Figure 5(c). This is consistent with the seen behaviors of water-restricted mice which first try to reach the water port and hydrate, after which their behaviors switch to exploring the maze or going back to the home state (Zhu et al., 2024; Ashwood et al., 2022). More details can be found in Appendix B.3.
@@
We further evaluate the quality of the recovered reward machine using a held-out set of unseen trajectories ($20$ test trajectories). The reward machine model and the product policy are learned from the remaining $180$ training trajectories. For each trajectory in the test set, we compute its log-likelihood under the product policy. This metric reflects how well the learned model captures the underlying trajectory distribution and generalizes beyond the training data. The results are shown in Table 4 and represent the average log-likelihood over the test dataset. To contextualize the results, we compare against three baselines: a uniformly random policy and two variants of Max Causal Entropy IRL Ziebart (2010). D-IRL employs a dense feature representation, assigning a distinct reward value to every MDP state. F-IRL uses a structured feature representation aligned with the environment’s labeling. Specifically, its feature vector is a one-hot encoding indicating whether the agent is at the home port ($\mathbf{h}$), at the water port ($\mathbf{w}$), or in any other intermediate state ($\mathbf{i}$). These results also quantitatively demonstrate the superiority of the learned reward machine in capturing the unseen behaviors of the mice.
Method D-IRL F-IRL Uniform LRM
Average log-likelihood 28.99-28.99 28.57-28.57 30.49-30.49 $\mathbf{-9.81}$
Table 4: Average log-likelihood performance of the learned reward machine on unseen trajectories, compared against three baseline methods.
6 Limitations and future work
The present framework assumes that the atomic propositions provided by the labeling function are exact, meaning no noise or mislabeling is present in the observed labels. While this assumption simplifies the analysis and ensures that the negative examples used in the SAT formulation are correct, it may be unrealistic in settings where the labeling function is derived from perception modules or noisy sensors. Future work could address this limitation by extending the framework to handle noise in the labeling function, for instance by integrating robust RM-learning techniques such as those in Parac et al. (2024), thereby broadening applicability to real-world domains.
Another limitation arises from the requirement to identify all negative examples up to depth $l=|\mathcal{S}|u_{\max}^{2}$ for the theoretical results to apply (see Theorem 1). Although this bound is polynomial in the size of the MDP and the maximum number of RM nodes, the number of distinct negative examples —and thus the number of clauses in the SAT formulation— may grow exponentially with $l$ in the worst case. For large problems, this can lead to SAT instances of prohibitive size and render the approach computationally intractable. On the other hand, in practice, we were able to identify the reward machine model with a much smaller depth in all our examples. Therefore, the existence of a tighter sufficient depth bound is an open problem, which we will investigate further in the future. We are also interested in developing smart prefix-selection algorithms that incorporate a verification loop to check, on the fly, whether the selected negative examples are sufficient.
@@ 7 Conclusion
Acknowledgments: This work is supported in part by ONR CLEVR-AI MURI (#N00014- 21-1-2431).
Appendices
@@ A.1 Proof of Proposition 1
More precisely, $\delta_{\textbf{u}}^{*,\mathrm{learned}}(u_{I},\sigma)=\delta_{\textbf{u}}^{*,\mathrm{learned}}(u_{I},\sigma^{\prime})$. Due to the contrapositive of Lemma 1, we have a contradiction. Similarly, due to the full specification condition of Section 3.1 and (10), the support of the two prefix policies will be the same by construction. ∎
@@ B.1 Tabular BlockWorld MDP
2Construct product MDP $\mathcal{M}_{\mathrm{Prod}}=\bigl(\mathcal{S}^{\prime},\;\mathcal{A}^{\prime},\;\mathcal{P}^{\prime},\;\mu_{0}^{\prime},\;\gamma^{\prime},\;r^{\prime}\bigr)$ as in Section $2.$
@@ 13 $\pi^{*}(a\mid s,u)\;\leftarrow\;\exp\!\bigl(Q_{k}(s,u,a)-V_{k}(s,u)\bigr)$
Input: MDP $\mathcal{M}$, reward machine $\mathcal{R}$, true product policy $\pi^{*}(a\mid s,u)$, # trajectories NN, trajectory length $H$
@@ 13    $s\leftarrow s^{\prime}$;
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ # SAT solutions
1000 20 58 0.05 24
3000 20 174 0.05 12
5000 20 305 0.05 8
10000 20 490 0.05 8
0.1M 20 2249 0.05 4
1M 20 7995 0.05 2
@@
$|\mathfrak{D}|$ $|\tau|$ $|\mathcal{E}^{-}|$ $\alpha$ # SAT solutions
200 20 26 0.05 32
500 20 52 0.05 8
1000 20 67 0.05 8
0.1M 20 1615 0.05 4
1M 20 5330 0.05 4
@@
B.2 2-Link Reacher Robotic Arm with Continuous State Space
@@ B.2.2 Discretizing the state space and sampling the policy
The MDP state space is generated by discretizing the shoulder angle $\theta_{1}$, the elbow angle $\theta_{2}$, and the angular velocities $\dot{\theta}_{1},\dot{\theta}_{2}$. The angles are discretized uniformly in $\pi,\pi]$ with a bin size of $10^{\circ}\approx 0.17$ radians. The angular velocities are discretized uniformly in [14,14][-14,14] with a bin size of $0.25$ rad/s. This leads to $\approx 17.4M$ states in the MDP.
@@ B.3 Real-world Mice Navigation
For recovering the reward function, we first need to construct the product policy. However, real-world data generally do not perfectly fit the assumed mathematical models, i.e., the max entropy assumption. Since most of the states in the MDP are unvisited, and many action are not sampled at most of the states, the product policy is predominantly sparse ($\sim 92\%$ of the action probabilities are 0). We preprocess the policy by clipping its minimum value to $0.05$. This allows us to use the results from [Shehab et al., 2024] to compute the reward function, as we require computing the $\log$ of the policy. The reward space is constrained to satisfy the feature-decomposition which the reward machine provides. However, there is generally no guarantee that we can find such a featurized reward function that can perfectly induce any product policy. With the shown reward in Figure 5(c), the norm difference between the induced policy and the true policy is $2.66$ (the norm difference with a uniformly random policy is $2.97$). By increasing $u_{\mathrm{max}}$ to $3$, one of the recovered reward machines is shown in Figure 10, which yields a slightly larger policy difference norm of $2.71$.
@@
Figure 10: Reward machine model considered in Appendix C.
Appendix C Policy Optimality is not Required
In this section, we illustrate with a toy example that the optimality of the prefix tree policy is not necessary to recover the reward machine model. We consider the MDP given by the BlockWorld of Section 5.2, and a simple 2-node reward machine model given by Figure 10. We do not specify a reward function here nor solve the forward MaxEnt RL problem before simulating the dataset, and instead specify an arbitrary stochastic policy at each node of the reward machine. With a dataset size $50K$ and trajectory length $20$, our method learns uniquely the reward machine model.
Appendix D Sensitivity to the Node Bound $u_{\mathrm{max}}$
To understand the effect of misspecifying the number of nodes on the resulting learned reward machine, we run an additional experiment in the Gridworld benchmark (see Section 5.1) while varying the bound on the number of nodes. We simulate a dataset of $2,500$ trajectories of length $10$ each. We then construct the negative example set using our finite sample results and solve for a reward machine model using MaxSAT. The recovered models are shown in Figure 11 below. These reward machine models still qualitatively capture the overall task even with the limited number of nodes. To quantitatively assess the quality of these models, we first learn a product policy for each using Algorithm 4 below. We then perform trajectory rollouts to compute the return of the policy consistent with each model. While the actions are sampled according to the corresponding learned policy of each model, the reward is calculated based on the traversed nodes in the ground-truth reward machine, given by Figure 1(b) [^2] For this experiment we set the reward for completing the task at $10$.. Results in Table 7 show a graceful degradation in performance as the number of reward machine nodes decreases. We also implemented a Max Entropy IRL baseline Ziebart et al. [2008], where a static reward is learned from the demonstration set.
(a) $u_{\mathrm{max}}=2$
(b) $u_{\mathrm{max}}=3$
Figure 11: Recovered reward machine models with varying node bounds.
Model Total #\# clauses #\# used clauses Rollouts $|\tau|$ Average Return per Rollout
$u_{\mathrm{max}}=4$ 1112 1112 10K 100 176.62
$u_{\mathrm{max}}=3$ 1112 902 10K 100 148.74
$u_{\mathrm{max}}=2$ 1112 589 10K 100 94.25
D-IRL - - 10K 100 17.12
F-IRL - - 10K 100 14.49
Table 7: Statistics for the varying bound experiment. (Total #\# clauses) is the number of negative examples from the finite dataset. (#\# used clauses) is the number of negative examples which MaxSAT included as hard constraints. (Rollouts) is the total number of rollouts. $(|\tau|)$ is the length of the trajectory. (Average Return per Rollout) is the sum of accumulated rewards in all the rollouts divided by the number of rollouts.
Input: Learned reward machine model $\widehat{\mathcal{G}}$, Ground-truth reward machine model $\mathcal{G}$, True prefix tree policy $\pi_{\mathrm{PTP}}$, Labeling function $L$, MDP $\mathcal{M}$, prefix length $d_{p}$
Output: Learned product policy $\widehat{\pi}$
1
2foreach node $u\in\widehat{\mathcal{G}}$ do
3   Compute $\text{Reach}(u)$ ;
 // we consider prefixes up to a fixed length of $d_{p}$
4
5 foreach $\sigma\in\mathrm{Paths}(u)$ do
6      Compute $\mathrm{Reach}(\sigma)$;
7
8    foreach $s\in\mathrm{Reach}(\sigma)$ do
9       foreach $a\in\mathcal{A}$ do
10          $\widehat{\pi}(a\mid s,u)\mathrel{+}=\pi_{\mathrm{PTP}}(a\mid s,\sigma)$
11
12
13
14Normalize $\widehat{\pi}$ and fill in zero rows with uniform policy;
return $\widehat{\pi}$
Algorithm 4 Construct Learned Product Policy
Appendix E Transferability Performance
In this section, we test the transferability of our recovered reward machines when deployed on a different environment with the same labeling function. In particular, the new environment is similar to the Tabular GridWorld MDP studied in Section 5.1, except that the room assignment is changed, as shown in Figure 12(a) below. The dynamics are kept the same. In order to test transferability, we first generate a trajectory dataset using the ground-truth optimal policy in the new environment. Then, for each reward machine model, given by a varied $u_{\mathrm{max}}$, we construct a product policy according to Algorithm 4 (using the prefix tree policy of the original environment) and solve for a reward function that will be optimized in the new environment. We then compute the log-likelihood of the trajectory dataset given the corresponding optimized policies. The results are shown in Figure 12(b), where (TRUE) is the log-likelihood of the ground-truth optimal policy in the new environment. It can be seen that the performance of the learned reward machine with $4$ nodes almost matches that of the ground-truth reward machine. The results also display a graceful degradation in performance with decreasing number of nodes.
Figure 12: (a): Modified labeled MDP. (b): Average log-likelihood of different models over 100 trajectories generated by the optimal policy in the new labeled MDP.